# Universal quantum control over Majorana zero modes

### v1

@@ II.2 Braiding operations via UQC passage
In this subsection, we construct the braiding operations between arbitrary pair of three MZMs, i.e., $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$, through a synthesis of the effective Hamiltonian Jin and Jing (2025d); James and Jerke (2007) and UQC theory Jin and Jing (2025a, b, c, d) (see Appendix A for details). In the regime where the mediator LD is under the largely detuned driving fields, the system dynamics is governed by the effective Hamiltonian Jin and Jing (2025d); James and Jerke (2007) as
$H_{\rm eff}(t)=-\frac{i}{2}\left[H_{I}(t),\int_{0}^{t}H_{I}(t_{1})dt_{1}\right].$ (8)
In the control Hamiltonian $H_{I}(t)$ in Eq. (7), the Rabi frequencies are assumed to be isotropic $\Omega_{1}(t)=\Omega_{2}(t)=\Omega_{3}(t)$ and the detunings are set as $\Delta_{j}(t)=\Delta_{0}-\Delta(t)/2$ and $\Delta_{k}(t)=\Delta_{0}+\Delta(t)/2$ with the scaling detunings $\Delta_{0}$ and $\Delta(t)$. Then under the large detuning condition $\Delta_{j}(t),\Delta_{k}(t),|\Delta_{j}(t)-\Delta_{l}(t)|,|\Delta_{k}(t)-\Delta_{l}(t)|\gg\Omega_{1}(t)$ with $l\neq j,k\in\{1,2,3\}$ and the near-resonant condition $\Delta_{j}(t)-\Delta_{k}(t)=\Delta(t)\sim\Omega_{1}(t)$, the effective Hamiltonian (8) results in a two-body interaction Hamiltonian:
@@
where the approximation holds under $\Delta_{0}\gg\Delta(t)/2$. In the double-rotated picture with respect to $H_{\rm MZM}=\Delta(t)(|\gamma_{1,k}\rangle\langle\gamma_{1,k}|-|\gamma_{1,j}\rangle\langle\gamma_{1,j}|)/2$, we have
$H_{\rm eff}^{\prime}(t)$ $=\frac{\Delta(t)}{2}(|\gamma_{1,j}\rangle\langle\gamma_{1,j}|-|\gamma_{1,k}\rangle\langle\gamma_{1,k}|)$ (11)
@@
whose level transition diagram is shown in Fig. 1(b). $\Delta(t)$ in Eq. (11) is directly controlled by the time-dependent detunings in laboratory, i.e., $\Delta_{j}(t)$ and $\Delta_{k}(t)$, rather than the extra driving fields in the previous protocol for effective Hamiltonian Jin and Jing (2025d).
Using the universal quantum control theory Jin and Jing (2025a, b, c, d) for an effective two-level system, the dynamics of the two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$ can be described in the ancillary representation spanned by
@@
$+e^{-if(\theta,\alpha)}|\mu_{2}(\theta,\alpha)\rangle\langle\mu_{2}(\theta_{0},\alpha_{0})|,$
where the subscript 0 indicates the initial boundary condition and the global phase $f(\theta,\alpha)$ is determined by
@@
To avoid the singularity of the parameters $\Delta(t)$ and $\Omega(t)$ in Eq. (13), one can take $\theta(t)$, $\alpha(t)$, and $f(\theta,\alpha)$ as independent variables. The conditions in Eq. (13) can then be expressed as
@@
Figure 2: Fidelity $\mathcal{F}(\theta)$ versus the parameter $\theta(t)$ for the states $|n\rangle$, $n=\gamma_{1,j},\gamma_{1,k},\gamma_{1,l},e$, during the braiding process with the initial population in MZM (a) $\gamma_{1,j}$ and (b) $\gamma_{1,k}$, respectively. The Kitaev-chain size is $N=100$. $\Delta_{j}(t)=\Delta_{0}-\Delta(t)/2$ and $\Delta_{k}(t)=\Delta_{0}+\Delta(t)/2$ where $\Delta(t)$ in Eq. (16a) and it is in the order of $\Delta(t)\sim 10^{-2}\Delta_{0}$. $\Delta_{l}=2\Delta_{0}$. Under the Rabi frequencies of $\Omega_{1}(t)=\Omega_{2}(t)=\Omega_{3}(t)$, which satisfy Eqs. (10) and (16c) with $\theta(t)=\pi t/(2T)$ ($T$ is the control period) and $f(\theta,\alpha)=0$. The Rabi frequency is in the order of $\Omega_{1}(t)\sim 10^{-2}\Delta_{0}$.
@@
When $f(\theta,\alpha)=0$, it is further reduced to
$U(\theta,\alpha)=-|\gamma_{1,k}\rangle\langle\gamma_{1,j}|+|\gamma_{1,j}\rangle\langle\gamma_{1,k}|=\exp\left(\frac{\pi}{4}\gamma_{1,j}\gamma_{1,k}\right),$ (18)
which is consistent with the braiding operation between the two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$ Ivanov (2001); Stone and Chung (2006); Read (2009); Sau et al. (2011); van Heck et al. (2012); Kraus et al. (2013).
Due to the transition diagram in the inset of Fig. 1(a), the performance of our braiding protocol can be evaluated by the fidelities for all the four relevant states: $\mathcal{F}(\theta)\equiv|\langle n|\psi(\theta)\rangle|^{2}$, $n=\gamma_{1,j},\gamma_{1,k},\gamma_{1,l},e$, where the wave function $|\psi(\theta)\rangle$ is obtained by the Schrödinger equation $i\partial|\psi(\theta)\rangle/\partial t=H_{\rm tot}(t)|\psi(\theta)\rangle$ with the total Hamiltonian $H_{\rm tot}(t)$ (4). Figures 2(a) and (b) present the numerical results of the fidelities starting from different initial MZMs. They are exact the same as the braiding process controlled by $U(\theta,\alpha)$ in Eq. (18), during which the states $|\gamma_{1,l}\rangle$ and $|e\rangle$ are effectively decoupled from the whole evolution. Particularly, in Fig. 2(a), the initial state $|\gamma_{1,j}\rangle$ evolves into an equal superposition of the states $|\gamma_{1,j}\rangle$ and $|\gamma_{1,k}\rangle$ when $\theta=\pi/4$ and then completely transfers to the state $|\gamma_{1,k}\rangle$ when $\theta=\pi/2$. When the system is initially populated on MZM $\gamma_{1,k}$, it can be faithfully transferred to the target MZM $\gamma_{1,j}$ when $\theta(T)=\pi/2$ as shown in Fig. 2(b). In the languages of UQC, Figures 2(a) and (b) simulate the nonadiabatic and transitionless passages $|\mu_{1}\rangle$ and $|\mu_{2}\rangle$ in Eq. (12), respectively.
Our protocol can be straightforwardly generalized from Eq. (7) of $3$ MZMs to the case with $M$ MZMs, i.e., the control Hamiltonian can be scaled to
@@
The preceding construction of braiding operation can be implemented between any selected pair of MZMs, e.g., the $j$th and $k$th MZMs, $j\neq k$. It is found that under the isotropic condition for Rabi frequencies $\Omega_{l}(t)=\Omega_{1}(t)$ with $1\leq l\leq M$, the large detuning condition $\Delta_{j}(t),\Delta_{k}(t),|\Delta_{j}(t)-\Delta_{l\neq j,k}(t)|,|\Delta_{k}(t)-\Delta_{l\neq j,k}(t)|\gg\Omega_{1}(t)$, and the near-resonant condition $\Delta_{j}(t)-\Delta_{k}(t)\sim\Omega_{1}(t)$, one can obtain the same two-body effective Hamiltonian in Eq. (9) by Eq. (8). Thus, our braiding protocol is scale free.
@@ II.3 Suppress systematic errors in local defect
While the braiding operations of MZMs remain robust against local perturbations in the other Majorana fermions, this topological protection does not cover the systematic errors about the mediator LD. As indicated by Eq. (47), the unavoidable systematic errors associated with LD can cause undesirable transitions among the universal passages. Incorporating an error correction mechanism in e.g., Eq. (53) to our protocol provides a reliable method to mitigate these systematic errors during the braiding operations of MZMs.
From a pedagogical perspective, the full Hamiltonian $H_{\rm rot}(t)$ in Eq. (4) can be perturbed by: (a) the fluctuation in the splitting frequency of the local defect, i.e.,
@@
or (b) the Rabi-frequency fluctuation in the driving field for one of the transitions $|e\rangle\leftrightarrow|\gamma_{1,1}\rangle$, i.e.,
@@
where the factor $\lambda$ scales the variation rapidity of $f(\theta)$ relative to $\theta(t)$. $\lambda=0$ represents the parallel-transport condition with a static global phase Jin and Jing (2025c) and $\lambda\gg 1$ describes a rapid-varying global-phase that can suppress the systematic errors in the relevant nonadiabatic passage.
@@
Figure 3: Fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the perturbative coefficient $\epsilon$, controlled by the full Hamiltonian $H_{\rm tot}(t)$ (4) with (a) deviated transition frequency $\omega$ in Eq. (20) and (b) deviated Rabi-frequency $\Omega_{1}(t)$ in Eq. (21), under the global phase $f(\theta)$ in Eq. (22) with various coefficients $\lambda$. The other parameters are the same as Fig. 2(a).
@@
Figures 3(a) and (b) demonstrate the fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the dimensionless error magnitude $\epsilon$, in the presence of the systematic errors in Eqs. (20) and (21), respectively. One can find that the braiding operations incorporating with our correction mechanism ($\lambda>1$) exhibits superior performances in comparison to those with no correction mechanism ($\lambda=0$). In Fig. 3(a) for $\lambda=0$, the fidelity decreases to $\mathcal{F}=0.924$ when $\epsilon=-0.05$ and $\mathcal{F}=0.807$ when $\epsilon=0.05$. Under our correction mechanism, the fidelity admits a lower bound as $\mathcal{F}>0.963$ when $\lambda=2$ and $\mathcal{F}>0.995$ when $\lambda\geq 5$, in the whole range of $\epsilon\in[-0.05,0.05]$. As shown in Fig. 3(b), the braiding operation with our correction mechanism maintains a stronger robustness against the deviation in the Rabi-frequency of a local driving field. In particular, when $\lambda=0$, the fidelity is about $\mathcal{F}=0.962$ when $\epsilon=-0.1$ and $\mathcal{F}=0.968$ when $\epsilon=0.1$. The adverse effects from the local errors can be suppressed so that $\mathcal{F}>0.980$ when $\lambda=2$, and $\mathcal{F}>0.998$ when $\lambda\geq 5$.

### v2

@@ II.2 Braiding operations via UQC passage
In this subsection, we construct the braiding operations between an arbitrary pair in three MZMs, i.e., $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$, through a synthesis of the effective Hamiltonian Jin2025Rydberg; James2007Effective and UQC theory Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr; Jin2025Rydberg (see Appendix A for details). When the mediator LD is under the largely detuned driving fields, the system dynamics is governed by the effective Hamiltonian Jin2025Rydberg; James2007Effective
$H_{\rm eff}(t)=-\frac{i}{2}\left[H_{I}(t),\int_{0}^{t}H_{I}(s)ds\right].$ (8)
For simplicity, the Rabi frequencies in the control Hamiltonian (7) are set to be isotropic, i.e., $\Omega_{1}(t)=\Omega_{2}(t)=\Omega_{3}(t)$. The detunings are parameterized as $\Delta_{j}(t)=\Delta_{0}-\Delta(t)/2$ and $\Delta_{k}(t)=\Delta_{0}+\Delta(t)/2$ with the scaling detunings $\Delta_{0}$ and $\Delta(t)$. Then under the large detuning condition $\Delta_{j}(t),\Delta_{k}(t),|\Delta_{j}(t)-\Delta_{l}(t)|,|\Delta_{k}(t)-\Delta_{l}(t)|\gg\Omega_{1}(t)$ with $l\neq j,k\in\{1,2,3\}$ and the near-resonant condition $\Delta_{k}(t)-\Delta_{j}(t)=\Delta(t)\sim\Omega_{1}(t)$, the effective Hamiltonian (8) results in a two-body interaction Hamiltonian:
@@
where the approximation holds under $\Delta_{0}\gg\Delta(t)$. In the derivation of Eq. (9), the first exponential terms in Eq. (7) have been neglected under the rotating wave approximation. As for the products by the second exponential terms, only those proportional to $\exp[i\int_{0}^{t}\Delta(s)ds]$ are slowly oscillating with time and then can be retained. In the double-rotated picture with respect to $H_{\rm MZM}=\Delta(t)(|\gamma_{1,k}\rangle\langle\gamma_{1,k}|-|\gamma_{1,j}\rangle\langle\gamma_{1,j}|)/2$, we have
$H_{\rm eff}^{\prime}(t)$ $=\frac{\Delta(t)}{2}\left(|\gamma_{1,j}\rangle\langle\gamma_{1,j}|-|\gamma_{1,k}\rangle\langle\gamma_{1,k}|\right)$ (11)
@@
whose level transition diagram is shown in Fig. 1(b). $\Delta(t)$ in Eq. (11) is directly controlled by the time-dependent detunings in laboratory, i.e., $\Delta_{j}(t)$ and $\Delta_{k}(t)$, rather than the extra driving fields in the previous protocol for effective Hamiltonian Jin2025Rydberg.
Using the universal quantum control theory Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr; Jin2025Rydberg for an effective two-level system, the dynamics of the two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$ can be described in the ancillary representation spanned by
@@
$+e^{-if(\theta,\alpha)}|\mu_{2}(\theta,\alpha)\rangle\langle\mu_{2}(\theta_{0},\alpha_{0})|.$
Here the subscript 0 indicates the initial boundary condition and the global phase $f(\theta,\alpha)$ is determined by
@@
The second equivalence in the last equation used the second condition in Eq. (13). To avoid the singularity of the parameters $\Delta(t)$ and $\Omega(t)$ in Eq. (13), one can regard $\theta(t)$, $\alpha(t)$, and $f(\theta,\alpha)$ in Eq. (15) as independent variables. The conditions in Eq. (13) can then be expressed as
@@
Figure 2: Fidelity $\mathcal{F}(\theta)$ for the states $|n\rangle$, $n=\gamma_{1,j},\gamma_{1,k},\gamma_{1,l},e$, as a function of the parameter $\theta(t)$ during the braiding process with the initial population in MZMs (a) $\gamma_{1,j}$ and (b) $\gamma_{1,k}$. $\Delta_{j}(t)=\Delta_{0}-\Delta(t)/2$ and $\Delta_{k}(t)=\Delta_{0}+\Delta(t)/2$ where $\Delta(t)$ in Eq. (16a) and it is in the order of $\Delta(t)\sim 10^{-2}\Delta_{0}$. $\Delta_{l}=2\Delta_{0}$. Under $\Omega_{1}(t)=\Omega_{2}(t)=\Omega_{3}(t)$, the effective Rabi frequency $\Omega(t)$ satisfies Eqs. (10) and (16c) with $\theta(t)=\pi t/(2T)$ ($T$ is the control period) and $f(\theta,\alpha)=0$. The Rabi frequency is in the order of $\Omega_{1}(t)\sim 10^{-2}\Delta_{0}$.
@@
When $f(\theta,\alpha)=0$, it is further reduced to a typical braiding operation between the two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$ Ivanov2001NonAbelian; Stone2006Fusion; Read2009NonAbelian; Sau2011Controlling; vanHeck2012Coulomb; Kraus2013Braiding:
$U(\theta,\alpha)=-|\gamma_{1,k}\rangle\langle\gamma_{1,j}|+|\gamma_{1,j}\rangle\langle\gamma_{1,k}|=\exp\left(\frac{\pi}{4}\gamma_{1,j}\gamma_{1,k}\right).$ (18)
Due to the transition diagram in the inset of Fig. 1(a), the performance of our braiding protocol can be evaluated by the fidelities for all the four relevant states: $\mathcal{F}(\theta)\equiv|\langle n|\psi(\theta)\rangle|^{2}$, $n=\gamma_{1,j},\gamma_{1,k},\gamma_{1,l},e$, where the wave function $|\psi(\theta)\rangle$ is directly calculated by the Schrödinger equation $i\partial_{t}|\psi(\theta)\rangle=H_{\rm tot}(t)|\psi(\theta)\rangle$ with the total Hamiltonian (4). Figures 2(a) and (b) present the numerical results of the fidelities starting from different MZMs. They are exactly the same as the braiding process controlled by $U(\theta,\alpha)$ in Eq. (18), during which the states $|\gamma_{1,l}\rangle$ and $|e\rangle$ are effectively decoupled from the whole evolution. Particularly, in Fig. 2(a), the initial state $|\gamma_{1,j}\rangle$ evolves into an equal superposition of the states $|\gamma_{1,j}\rangle$ and $|\gamma_{1,k}\rangle$ when $\theta=\pi/4$ and then completely transfers to the state $|\gamma_{1,k}\rangle$ when $\theta=\pi/2$. When the system is initially populated on MZM $\gamma_{1,k}$, it can be faithfully transferred to the target MZM $\gamma_{1,j}$ when $\theta(T)=\pi/2$ as shown in Fig. 2(b). In the languages of UQC, Figures 2(a) and (b) simulate the nonadiabatic and transitionless passages $|\mu_{1}\rangle$ and $|\mu_{2}\rangle$ in Eq. (12), respectively.
Our protocol can be straightforwardly generalized from $3$ MZMs to $M$ MZMs. In this case, the control Hamiltonian (7) will be rescaled to
@@
To modulate the hopping and pairing interaction between LD and the end sites of $M$ Kitaev chains, a roughly linear-scaling experimental overhead has to be paid to control the eigenfrequencies of the fermions and their neighboring couplings in the Kitaev chains, including $\mu$, $J$, and $g$. Then the preceding construction of braiding operation can be implemented between any selected pair of MZMs, e.g., the $j$th and $k$th MZMs, $j\neq k$. It is found that under the isotropic condition for Rabi frequencies $\Omega_{k}(t)=\Omega_{1}(t)$ with $1\leq k\leq M$, the large detuning condition $\Delta_{j}(t),\Delta_{k}(t),|\Delta_{j}(t)-\Delta_{l\neq j,k}(t)|,|\Delta_{k}(t)-\Delta_{l\neq j,k}(t)|\gg\Omega_{1}(t)$, and the near-resonant condition $|\Delta_{j}(t)-\Delta_{k}(t)|\sim\Omega_{1}(t)$, one can again obtain the same two-body effective Hamiltonian as in Eq. (9) by Eq. (8).
@@ II.3 Suppress systematic errors in local defect
While the braiding operations of MZMs are resilient to the local perturbations in the other Majorana fermions, this topological protection does not cover the systematic errors about the mediator LD. As indicated by Eq. (52), the unavoidable systematic errors associated with LD can cause undesirable transitions among the universal passages. A reliable mitigating method is incorporating an error correction mechanism through e.g., Eq. (58) to our protocol during the braiding operations of MZMs.
From a pedagogical perspective, the full Hamiltonian $H_{\rm rot}(t)$ in Eq. (4) can be perturbed by: (i) the fluctuation in the splitting frequency of the local defect, i.e.,
@@
or (ii) the Rabi-frequency fluctuation in the driving field for one of the transitions $|e\rangle\leftrightarrow|\gamma_{1,1}\rangle$, i.e.,
@@
where $\lambda$ scales the variation rate of $f(\theta)$ with respect to $\theta(t)$. $\lambda=0$ represents the parallel-transport condition with a static global phase Jin2025ErrCorr and $|\lambda|\gg 1$ describes a rapid-varying global-phase that can suppress the systematic errors in the relevant nonadiabatic passage. Indicated by Eqs. (13) and (15), the global phase is uniquely determined by $\Delta(t)$ and $\Omega(t)$, through the parameters $\theta(t)$ and $\alpha(t)$ in universal passage. Experimentally, the rapid modulation of the global phase can be achieved by the gate voltages of the superconductor circuit embedded in the hybrid semiconductor-superconductor systems Dvir2023Realization; Shi2024Probing; Bordin2025Enhanced; Microsoft2025Interferometric; Nitsch2025Poor, where the superconductor circuit is adiabatically eliminated to effectively realize the tunable hopping and pairing couplings between QDs as described by Eq. (6).
@@
Figure 3: Fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the perturbative coefficient $\epsilon$, controlled by the full Hamiltonian $H_{\rm tot}(t)$ (4) with (a) deviated transition frequency $\omega$ in Eq. (20) and (b) deviated Rabi frequency $\Omega_{1}(t)$ in Eq. (21), under the global phase $f(\theta)$ in Eq. (22) with various coefficients $\lambda$. The other parameters are the same as Fig. 2(a).
@@
Figures 3(a) and (b) demonstrate the fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the dimensionless error magnitude $\epsilon$, in the presence of the systematic errors in Eqs. (20) and (21), respectively. One can find that the braiding operations incorporating with our correction mechanism ($|\lambda|>1$) exhibits superior performances in comparison to those with no correction mechanism ($\lambda=0$). In Fig. 3(a) for $\lambda=0$, the fidelity declines to $\mathcal{F}=0.924$ when $\epsilon=-0.05$ and $\mathcal{F}=0.807$ when $\epsilon=0.05$. Under our correction mechanism, the fidelity admits a lower bound as $\mathcal{F}>0.963$ when $\lambda=2$ and $\mathcal{F}>0.995$ when $\lambda\geq 5$, in the whole range of $\epsilon\in[-0.05,0.05]$. As shown in Fig. 3(b), the braiding operation with our correction mechanism exhibits a more significant robustness against the deviation in Rabi frequency of a local driving field. In particular, when $\lambda=0$, the fidelity is about $\mathcal{F}=0.962$ when $\epsilon=-0.1$ and $\mathcal{F}=0.968$ when $\epsilon=0.1$. The adverse effects from the local errors can be suppressed so that $\mathcal{F}>0.980$ when $\lambda=2$, and $\mathcal{F}>0.998$ when $\lambda\geq 5$.
II.4 Suppress quasiparticle poisoning
The fermion parity of the subspace of Majorana zero modes is conserved with the systematic errors on local defect as discussed in Sec. II.3. However, in practical topological quantum computation, the fermion parity can be broken by quasiparticle poisoning Flsensberg2010Tunneling; Leijnse2011Scheme; Rainis2012Majorana; Karzig2021Quasiparticle. It occurs when an undesired Majorana fermion participates the braiding operation and typically arises from the nonideal conditions of the system, including environmental noise, perturbations in system parameters, and thermal excitations Karzig2021Quasiparticle.
We consider the following Hamiltonian to simulate the effect from the quasiparticle poisoning Knapp2016Nature:
$\tilde{H}_{\rm tot}(t)=H_{\rm eff}(t)+H_{\rm err}(t),$ (23)
where $H_{\rm eff}(t)$ is given by Eq. (9) with $\Omega(t)$ the effective Rabi-frequency of the transition between the interested states $|\gamma_{1,j}\rangle$ and $|\gamma_{1,k}\rangle$, and the error Hamiltonian $H_{\rm err}(t)$ is given by
$H_{\rm err}(t)=\epsilon\Omega(t)|\gamma_{1,j}\rangle\langle\gamma_{1,l}|+{\rm H.c.},$ (24)
with $\epsilon$ the leakage magnitude out of the topological subspace spanned by $|\gamma_{1,j}\rangle$ and $|\gamma_{1,k}\rangle$. The quasiparticle poisoning described by Eq. (24) can arise from the unwanted hopping and pairing interactions between the fermions at the ends of the $j$th and $l$th Kitaev chains, i.e., $c_{1,j}$ and $c_{1,l}$.
Figure 4: Fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the perturbative coefficient $\epsilon$, controlled by the nonideal Hamiltonian $\tilde{H}_{\rm tot}(t)$ in Eq. (23). The other parameters are the same as Fig. 2(a).
In Fig. 4, we demonstrate the fidelity $\mathcal{F}[\theta(T)=\pi/2]$ about the target state $|\gamma_{1,k}\rangle$ versus the dimensionless error magnitude $\epsilon$, under the nonideal Hamiltonian (23). One can find that our correction mechanism in Eq. (22) is still applicable to suppress the effect from the quasiparticle poisoning. In particular, for $\lambda=0$, the fidelity declines to $\mathcal{F}[\theta(T)=\pi/2]=0.967$ when $\epsilon=\pm 0.2$. When $\lambda=5$, the fidelity is enhanced to $\mathcal{F}[\theta(T)=\pi/2]\geq 0.973$ in the whole range of $\epsilon\in[-0.2,0.2]$; and when $\lambda=10$, the lower-bound of the fidelity is over $0.984$ within the range of $\epsilon\in[-0.2,0.2]$.
Moreover, our protocol can be alternatively combined with the measurement-based error correction Knapp2016Nature to mitigate the quasiparticle poisoning. Particularly, the projection operator Knapp2016Nature for the $j$th and $k$th MZMs in our model can be defined as
$\Pi_{\pm}=\frac{1\pm i\gamma_{1,j}\gamma_{1,k}}{2}.$ (25)
In the presence of the leakage out of the topological subspace due to the unwanted crosstalk similar to Eq. (24), the ideal braid operation $U(\theta,\alpha)$ in Eq. (18) becomes
$U(\theta,\alpha)\rightarrow U_{\epsilon}=e^{\frac{\pi}{4}\gamma_{1,j}\gamma_{1,k}+\epsilon\gamma_{1,j}\gamma_{1,l}}$ (26)
$\approx e^{\frac{\pi}{4}\gamma_{1,j}\gamma_{1,k}}e^{\epsilon\gamma_{1,j}\gamma_{1,l}}e^{-\frac{1}{2}[\frac{\pi}{4}\gamma_{1,j}\gamma_{1,k},\epsilon\gamma_{1,j}\gamma_{1,l}]},$
$\approx U(\theta,\alpha)\left(1+\epsilon\gamma_{1,j}\gamma_{1,l}-\frac{\pi\epsilon}{8}[\gamma_{1,j}\gamma_{1,k},\gamma_{1,j}\gamma_{1,l}]\right)+\mathcal{O}(\epsilon^{2}),$
$=U(\theta,\alpha)\left(1+\epsilon\gamma_{1,j}\gamma_{1,l}-\frac{\pi\epsilon}{4}\gamma_{1,k}\gamma_{1,l}\right)+\mathcal{O}(\epsilon^{2}).$
Here we have applied the Baker-Campbell-Hausdorff formula with a small perturbation $\epsilon\ll 1$. Assume that the initial state of the system $|\psi_{i}\rangle$ satisfies the parity condition $i\gamma_{1,j}\gamma_{1,k}|\psi_{i}\rangle=1$. Then under the combination strategy, the final state justifies that the measurement-based error correction $\Pi_{+}$ in Eq. (25) can be used to suppress the leakage out of the topological subspace spanned by $\gamma_{1,j}$ and $\gamma_{1,k}$ to the first order:
$|\psi_{f}\rangle=\Pi_{+}U_{\epsilon}|\psi_{i}\rangle=\Pi_{+}U(\theta,\alpha)\Big(1+\epsilon\gamma_{1,j}\gamma_{1,l}$ (27)
$-\frac{\pi\epsilon}{4}\gamma_{1,k}\gamma_{1,l}\Big)|\psi_{i}\rangle+\mathcal{O}(\epsilon^{2})=U(\theta,\alpha)|\psi_{i}\rangle+\epsilon\Big[U(\theta,\alpha)$
$\times\Pi_{+}\gamma_{1,j}\gamma_{1,l}-\frac{\pi}{4}U(\theta,\alpha)\Pi_{+}\gamma_{1,k}\gamma_{1,l}\Big]|\psi_{i}\rangle+\mathcal{O}(\epsilon^{2})$
$=U(\theta,\alpha)|\psi_{i}\rangle+\epsilon\Big[U(\theta,\alpha)\gamma_{1,j}\gamma_{1,l}\Pi_{-}-\frac{\pi}{4}U(\theta,\alpha)$
$\times\gamma_{1,k}\gamma_{1,l}\Pi_{-}\Big]|\psi_{i}\rangle+\mathcal{O}(\epsilon^{2})=U(\theta,\alpha)|\psi_{i}\rangle+\mathcal{O}(\epsilon^{2}).$
In this derivation, we have used the anti-commutation relations $\{\gamma_{1,j}\gamma_{1,k},\gamma_{1,j}\gamma_{1,l}\}=0$ and $\{\gamma_{1,j}\gamma_{1,k},\gamma_{1,k}\gamma_{1,l}\}=0$, and the result $\Pi_{-}|\psi_{i}\rangle=0$.